how can we find the minimum length of a line ,which lies between the axes and passes through the point (a,b)?

Grade:12th pass

1 Answers

APURV GOEL
12 years ago

The minimum length of line will be when it makes 135 degree angle with +ve X-axis.

by y = mx + c

we have b = -a + c    as tan135 = -1

or c=a+b

So equation of the line is x + y = a + b

Please approve.